# Transfer-Based Strategies for Multi-Target Quantum Optimization

### v1

@@ Abstract
Index Terms:
multi-target optimization, transfer learning, quantum computing, parameterized quantum circuits
@@ I Introduction
Quantum computers are expected to solve hard computational problems significantly faster than their classical counterparts, based on superposition and entanglement properties. This advantage is demonstrated in many fields, from combinatorial optimization [1, 2], cryptography [3] to physical simulations [4]. Among these, quantum optimization has emerged as one of the most practical near-term applications. Quantum optimization promises to accelerate complex targets commonly encountered in classical optimization, such as processing large-scale input [2]. For instance, Quantum Approximate Optimization Algorithms (QAOA) are actively being developed to address combinatorial optimization problems [5, 1], while the Variational Quantum Eigensolver (VQE) is widely used for approximating eigenvalues in quantum chemistry [6, 7]. Moreover, Quantum Neural Networks (QNNs) are increasingly employed to tackle optimization targets within machine learning frameworks [8, 9, 10].
The multi-target quantum optimization (MTQO) problem is proposed in [11], which aims to optimize multiple objectives simultaneously, rather than a single objective as in previous works, which aligns with the classical formulation of multi-objective /target optimization [12]. Consider a set of $K$ optimization targets denoted as $T_{1},T_{2},\dots,T_{K}$, all defined over the same search space $\mathcal{S}$. Each target $T_{k}$ is associated with its own cost function $\mathcal{C}(\bm{\theta}_{(k)}):\mathbb{R}^{m}\to[0,1]$ where $\bm{\theta}_{(k)}$ is the $m$-dimensional parameter vector. A target $T_{k}$ that has the corresponding optimal parameter ($\bm{\theta}^{*}_{(k)}$) is marked as “optimized”. The goal of multi-target optimization is to find a set of solutions:
$\{\bm{\theta}^{*}_{(1)},\bm{\theta}^{*}_{(2)},\dots,\bm{\theta}^{*}_{(K)}\}$ (1)
, where $\bm{\theta}^{*}_{(k)}$ is the optimal solution for target $T_{k}$, with a given trainable unitary operator $U(\bm{\theta})$. There are two stages of multi-target optimization: one (training) is the stage where $|B|\gg|A|$, knowledge is transferred from the optimized one to nearby neighbors as a flooding algorithm, which is conducted efficiently through a D-level tree. The second stage (inference) is that $|A|\gg|B|$, each target $y_{k^{\prime}}\in B$ can be also estimated from $A$, visualized as Figure 1 (a). The unoptimized (white dot) can learn from the nearest optimized target (black dot).
In this work, we will explain the theoretical foundations of MTQO, its structure in Section III, and use techniques like the warm-start method and parameter estimator to enhance the MTQO process in Section IV. We also propose some deep learning techniques as more advantageous strategies. Finally, the initial comparison resources are presented in Section V.
@@ II Related works
For the first issue, i.e., what and how to transfer, researchers have primarily focused on designing effective transfer mechanisms to share useful information among targets. Existing approaches can be referred to as multi-population-based methods. Multi-population-based methods evolve separate populations for each target and perform the transfer between populations based on evolutionary information such as search distributions. Ref [15] proposed the multi-task genetic algorithm (MTGA) to minimize bias among target optima and to enable transfer through dimension-wise random shuffling. However, its random shuffling does not explicitly transfer optimization to account for dimensional similarity. Ref [16] proposed a multi-population framework with a novel mutation operator designed to enhance transferability between populations. More recently, research  [17] observed that transferring target-specific parameter alone may be insufficient and proposed a transfer approach that also transfers meta-knowledge between targets. This meta-knowledge captures higher-level strategies on how to obtain high-quality solutions and is thus more generalizable across targets with varying degrees of similarity.
@@ III Quantum optimization algorithm
The partial derivative $\partial_{\theta_{j_{(x)}}}\mathcal{C}(\bm{\theta}^{(t)}_{(x)},x)$ can be achieved via the PSR technique [23]. In case the generator of the quantum gate has two eigenvalues, such as one-qubit rotation gate, this technique is known as two-term PSR:
@@
where $e_{j}$ is the $j^{\text{th}}$-unit vector. For each target $x$, the Algorithm. 1 take $n_{\text{iter}(x)}$ iterations; each iteration refers to $2m+1$ quantum evaluations (QE) for cost value $\mathcal{C}(\bm{\theta}^{(t)}_{(x)},x)$ and its gradient. Totally, finding $\bm{\theta}^{*}_{(x)}$ require $n_{\text{iter}(x)}(2m+1)$ QEs. Note that $n_{\text{iter}(x)}$ grow exponentially by $n$.
@@ IV Strategies for multi-target quantum optimization
In many practical scenarios, optimization must be repeated across a family of input states where these inputs often reside in a continuous or structured space. Furthermore, solving 3 multiple times could lead to large resources on quantum computers. Assuming that we have $K$ target in a same Hilbert space $\mathcal{S}$ that divided into two subsets: $A=\{x_{(k)}\}_{k=1}^{K}$ and $B=\{y_{(k)}\}_{k^{\prime}=1}^{K^{\prime}}$, where $A$ includes the optimized targets and $B$ includes the unoptimized targets. We transfer the knowledge from $A$ to $B$ such that $\sum_{k=1}^{K^{\prime}}n_{\text{iter}(y_{k^{\prime}})}$ reduce as much as possible. Specifically, given a new input $y_{k}\in B$, we define a distance metric $d(y_{k^{\prime}},x_{k})$ on $\mathcal{S}$, and identify the nearest previously seen target $x_{k}$ such that:
@@
The strategies based on this method can be arranged from simple to advanced. In most simple way, the target $k^{*}$ can be transferred to $x_{k^{\prime}}$ if we use $\bm{\theta}^{*}_{(x_{k}^{*})}$ as parameter initialization (warm-starting method). A bit more complicated, we can use an estimator to move the parameters closer to the optimal point. By using the above-defined metric, the unoptimized targets can also be organized for better optimization performance. Finally, we suggest that applying deep learning methods which was successful in the classical multi-target optimization may help us improve the performance of quantum analogue.
@@ IV-A Warm-starting method
If $P(\bm{\theta}^{(0)},y_{k^{\prime}})$ take $n_{\text{iter}(y_{k^{\prime}})}$ iterations to converge. Since $y_{k^{\prime}}\approx x_{k^{*}}$ under the defined metric $d$, we expect $\bm{\theta}^{*}_{(y_{k^{\prime}})}$ to be close to $\bm{\theta}^{*}_{(x_{k^{*}})}$, which leads to a reduction in number of optimization iteration. In the other hand, $P(\bm{\theta}^{*}_{(x^{*}_{k})},y_{k^{\prime}})$ consume only $n^{*}_{\text{iter}(y_{k^{\prime}})}\ll n_{\text{iter}(y_{k^{\prime}})}$ iterations. This approach can be viewed as a warm-starting optimization method [25]. This method’s procedure is as follows. First, we optimize $\bm{\theta}_{(x)}$ for a target $x$ using randomly initialized parameters to minimize the cost value. Second, after the base PQC has been optimized, we fine-tune (continue to optimize) $\bm{\theta}^{*}_{(x)}$ to $\bm{\theta}^{*}_{(y)}$ for a new target $y$. Finally, we evaluate the effectiveness by comparing the performance of the PQC trained with the warm-starting and the PQC trained from scratch with default initialized parameters.
Figure 2: The average of #Iteration and cost $C_{(x)}(\bm{\theta}^{*}_{(x)})$. Each subplot represents one qubit configuration. The blue curves indicate the cost values obtained after optimization, while the red curves show the corresponding average of #Iteration consumed to reach those cost values. Solid lines denote results with warm-staring, whereas dashed lines represent the baseline without warm-staring.
@@ IV-C Organizing targets as a multi-level tree
${C}^{(1)},{C}^{(2)},\dots,{C}^{(D)}$ (13)
where each $\mathcal{C}^{(d)}=\{x^{(d)},C^{(d-1)}_{1},C^{(d-1)}_{2},\dots,C^{(d-1)}_{|d|}\}$ include a centroid $x^{(d)}$ and the set of tree at lower level d1d-1. Conversely, trees at level $d$ are formed by merging clusters from level d1d-1. The lower levels capture finer-grained clusters and higher levels capture more abstract groupings. This structure allows analysis at multiple scales, facilitating both local and global interpretations of the data.
Figure 1 (b) presents the 2-level tree system where level 0 has two sub-trees. The knowledge can be transferred from the centroid to sub-trees, for example: knowledge from the $\bm{\theta}^{*}_{d=0}$ is transferred to $\bm{\theta}_{d=1}$, knowledge from the $\bm{\theta}^{*}_{d=1}$ are transferred to $\bm{\theta}_{d=2}$ and so on. The optimization process on tree can be imagined as Figure 1 (c), instead of starting from $\bm{\theta}^{(0)}$, $\bm{\theta}^{*}_{d=1}$ and $\bm{\theta}^{*}_{d=2}$ can be obtained from $\bm{\theta}^{*}_{d=0}$ and $\bm{\theta}^{*}_{d=1}$, respectively. For example, we could take a hundred iterations for achieving $\bm{\theta}^{*}_{d=1}$ but only a few iterations if we start from the nearest neighbor $\bm{\theta}^{*}_{d=0}$ instead of $\bm{\theta}^{(0)}$.
@@ IV-D Deep learning method
$U(\bm{\theta}_{k})=\bigotimes_{j=0}\left(U_{\{j\}}\left(\bm{\theta}_{k\{j\}}\right)\right)$ (14)
where $\bm{\theta}_{k\{j\}}$ is the parameter $k$ at layer $j$; we can share mid-layer parameter for $\bm{\theta}_{k+1}$. Meanwhile, meta‑learning methods enable fast adaptation: a meta‑learner network is trained across a distribution of $k^{\prime\prime}\ll K$ targets, producing initial PQC parameters that can be fine‑tuned in just a few gradient steps for each new target. Studies suggest that such meta‑strategies significantly improve robustness under noise [30, 31]. Together, multi-task learning and meta‑learning comprise powerful tools for scalable, sample‑efficient optimization in quantum systems deploying multi‑target optimization.
@@ V Experiments
$U(\bm{\theta})=ZXZ\prod_{i=0}^{n-1}\text{CRY}(\theta)_{i,\ (i+1)\%n}$ (15)
The number of parameters and circuit depth of $n$-qubit $CR_{Y_{\text{chain}}}+ZXZ$ are $3nL$ and $6L$, respectively. The problem $P$ is implemented by Python 3.10 and Pennylane 0.41.0, and the scalability is investigated by increasing $n$ and $L$ ($L=n$). The maximum #Iteration and $\tau$ are fixed at 200 and $10^{-3}$ for all cases, respectively. Higher number of qubits, harder to optimize; in the other words, $P$ will require more iterations to achieve convergence.
@@ V-A Effectiveness of transfer-based strategies
TABLE I: Comparison of in warm-starting method and parameter estimator versus the no acceleration technique.
@@
$n$ Cost (TL) Cost Avg.#Iter (TL) Avg.#Iter
2 0.0008 0.0008 42.41 63.80
3 0.0009 0.0009 55.92 67.05
4 0.0009 0.0009 71.25 85.95
5 0.0011 0.0015 130.44 161.01
6 0.0153 0.0188 199.00 199.00
7 0.0874 0.0940 199.00 199.00
@@
$n$ Cost (TL) Cost Avg.#Iter (TL) Avg.#Iter
2 0.0008 0.0009 34.48 65.63
3 0.0009 0.0009 49.05 67.25
4 0.0009 0.0009 69.57 85.96
5 0.0011 0.0012 135.84 160.50
6 0.0155 0.0185 197.04 197.31
7 0.0888 0.0939 196.98 197.45
Table I summarizes all comparison results between transfer-based strategies and non-transfer-based strategies approaches, for two distinct methods: the warm-starting and the parameter estimator method. For each setting, we report the average of minimum cost values and the average number of iterations (Avg.#Iter) required to reach convergence. The warm-starting method is benchmarked following Figure 1 (a), given set $A,B$ where $|A|\gg|B|$, for each $y\in B$, we optimize $y$ with and without using knowledge from $A$. On the other hand, for the parameter estimator, we start with $B$ and an empty set $A$; for each optimized target, we move it from $B$ to $A$ until $B$ becomes an empty set. The first target will be optimized without the parameter estimator, and the latter will be transferred from $5$ nearest targets.
@@
Figure 2 shows the detailed result from the warm-starting method from 2-qubit to 5-qubit. Across all $K$, our method consistently achieves lower or comparable cost values using fewer iterations. In the 2-qubit case, the cost with warm-starting converges rapidly to a stable lower value, while the required #Iterations is significantly reduced compared to the baseline. This benefit is particularly pronounced for smaller systems, where the prior knowledge transferred from related tasks yields substantial optimization efficiency. In the 3-qubit and 4-qubit cases, the #Iteration for the warm-starting is again consistently lower, while both approaches maintain similar cost levels. For 5-qubit, as system complexity increases, the benefits of warm-starting become more clear where both cost and #Iteration are better. However, even in these cases, warm-starting demonstrates a trend toward requiring fewer iterations to reach comparable or slightly better performance, indicating its scalability. In higher-dimensional systems (6 to 8 qubits, Figure 3, right panel), the advantage of warm-starting becomes more pronounced. The dashed lines lie consistently below the solid lines, indicating lower final cost values even in more complex quantum systems. Furthermore, the cost values obtained via transfer are relatively stable as $K$ increases, in contrast to the baseline, which exhibits higher variability. This highlights the robustness of warm-starting in high-dimensional optimization landscapes.

### v2

@@ I Introduction
Quantum computers are expected to solve complex computational problems significantly faster than their classical counterparts, leveraging the properties of superposition and entanglement. This advantage is demonstrated in many fields, from combinatorial optimization [1, 2], cryptography [3] to physical simulations [4]. Among these, quantum optimization has emerged as one of the most practical near-term applications. Quantum optimization promises to accelerate complex targets commonly encountered in classical optimization, such as processing large-scale input [2]. For instance, Quantum Approximate Optimization Algorithms (QAOA) are actively being developed to address combinatorial optimization problems [5, 1], while the Variational Quantum Eigensolver (VQE) is widely used for approximating eigenvalues in quantum chemistry [6, 7]. Moreover, Quantum Neural Networks (QNNs) are increasingly employed to tackle optimization targets within machine learning frameworks [8, 9, 10].
The multi-target quantum optimization (MTQO) problem is proposed in [11], which aims to optimize multiple objectives simultaneously, rather than a single objective as in previous works, which aligns with the classical formulation of multi-objective (multi-target) optimization [12]. Consider a set of $K$ optimization targets denoted as $T_{1},T_{2},\dots,T_{K}$, all defined over the same search space $\mathcal{S}$. Each target $T_{k}$ is associated with its own cost function $\mathcal{C}(\bm{\theta}_{(k)}):\mathbb{R}^{m}\to[0,1]$ where $\bm{\theta}_{(k)}$ is the $m$-dimensional parameter vector. A target $T_{k}$ that has the corresponding optimal parameter ($\bm{\theta}^{*}_{(k)}$) is marked as “optimized”. The goal of multi-target optimization is to find a set of solutions:
$\{\bm{\theta}^{*}_{(1)},\bm{\theta}^{*}_{(2)},\dots,\bm{\theta}^{*}_{(K)}\},$ (1)
with $\bm{\theta}^{*}_{(k)}$ is the optimal solution for target $T_{k}$, with a given trainable unitary operator $U(\bm{\theta})$. There are two stages of multi-target optimization: one (training) is the stage with $|B|\gg|A|$, where $A$ includes the optimized targets and $B$ includes the unoptimized targets. The knowledge is transferred from the optimized one to nearby neighbors as a flooding algorithm, which is conducted efficiently through a D-level tree. The second stage (inference) is that $|A|\gg|B|$, each target $y_{k^{\prime}}\in B$ can be also estimated from $A$, visualized as Figure 1 (a). The unoptimized (white dot) can learn from the nearest optimized target (black dot).
In this work, we will explain the theoretical foundations of MTQO, its structure in Section III, and use techniques like the warm-start method and parameter estimator to enhance the MTQO process in Section IV. We also propose some deep learning techniques as more advantageous strategies. Finally, the initial comparative results are presented in Section V.
@@ II Related works
For the first issue, i.e., what and how to transfer, researchers have primarily focused on designing effective transfer mechanisms to share useful information among targets. Existing approaches can be referred to as multi-population-based methods. Multi-population-based methods evolve separate populations for each target and perform the transfer between populations based on evolutionary information, such as search distributions. Ref [15] proposed the multi-task genetic algorithm (MTGA) to minimize bias among target optima and to enable transfer through dimension-wise random shuffling. However, its random shuffling does not explicitly transfer optimization to account for dimensional similarity. Ref [16] proposed a multi-population framework with a novel mutation operator designed to enhance transferability between populations. More recently, research  [17] observed that transferring target-specific parameters alone may be insufficient and proposed a transfer approach that also transfers meta-knowledge between targets. This meta-knowledge captures higher-level strategies on how to obtain high-quality solutions and is thus more generalizable across targets with varying degrees of similarity.
@@ III Quantum optimization algorithm
The partial derivative $\partial_{\theta_{j_{(x)}}}\mathcal{C}(\bm{\theta}^{(t)}_{(x)},x)$ can be achieved via the PSR technique [23]. In case the generator of the quantum gate has two eigenvalues, such as one-qubit rotation gates, this technique is known as two-term PSR:
@@
$e_{j}$ is the $j^{\text{th}}$-unit vector. For each target $x$, Algorithm 1 takes $n_{\text{iter}(x)}$ iterations; each iteration refers to $2m+1$ quantum evaluations (QE) for cost value $\mathcal{C}(\bm{\theta}^{(t)}_{(x)},x)$ and its gradient. Totally, finding $\bm{\theta}^{*}_{(x)}$ require $n_{\text{iter}(x)}(2m+1)$ QEs. Note that $n_{\text{iter}(x)}$ grow exponentially by $n$.
@@ IV Strategies for multi-target quantum optimization
In many practical scenarios, optimization must be repeated across a family of input states where these inputs often reside in a continuous or structured space. Furthermore, solving (3) multiple times could lead to large resources on quantum computers. Assuming that we have $K$ target in a same Hilbert space $\mathcal{S}$ that divided into two subsets: $A=\{x_{(k)}\}_{k=1}^{K}$ and $B=\{y_{(k)}\}_{k^{\prime}=1}^{K^{\prime}}$, where $A$ includes the optimized targets and $B$ includes the unoptimized targets. We transfer the knowledge from $A$ to $B$ such that $\sum_{k=1}^{K^{\prime}}n_{\text{iter}(y_{k^{\prime}})}$ reduce as much as possible. Specifically, given a new input $y_{k}\in B$, we define a distance metric $d(y_{k^{\prime}},x_{k})$ on $\mathcal{S}$, and identify the nearest previously seen target $x_{k}$ such that:
@@
The strategies based on this method can be arranged from simple to advanced. In most simple way, the target $k^{*}$ can be transferred to $x_{k^{\prime}}$ if we use $\bm{\theta}^{*}_{(x_{k}^{*})}$ as parameter initialization (warm-starting method). A bit more complicated, we can use an estimator to move the parameters closer to the optimal point. By using the above-defined metric, the unoptimized targets can also be organized for better optimization performance. Finally, we suggest that applying deep learning methods, which were successful in the classical multi-target optimization, may help us improve the performance of the quantum analogue.
@@ IV-A Warm-starting method
If $P(\bm{\theta}^{(0)},y_{k^{\prime}})$ take $n_{\text{iter}(y_{k^{\prime}})}$ iterations to converge. Since $y_{k^{\prime}}\approx x_{k^{*}}$ under the defined metric $d$, we expect $\bm{\theta}^{*}_{(y_{k^{\prime}})}$ to be close to $\bm{\theta}^{*}_{(x_{k^{*}})}$, which leads to a reduction in number of optimization iteration. On the other hand, $P(\bm{\theta}^{*}_{(x^{*}_{k})},y_{k^{\prime}})$ consume only $n^{*}_{\text{iter}(y_{k^{\prime}})}\ll n_{\text{iter}(y_{k^{\prime}})}$ iterations. This approach can be viewed as a warm-starting optimization method [25]. The procedure of this method is as follows. First, we optimize $\bm{\theta}_{(x)}$ for a target $x$ using randomly initialized parameters to minimize the cost value. Second, after the base PQC has been optimized, we fine-tune (continue to optimize) $\bm{\theta}^{*}_{(x)}$ to $\bm{\theta}^{*}_{(y)}$ for a new target $y$. Finally, we evaluate the effectiveness by comparing the performance of the PQC trained with the warm-starting and the PQC trained from scratch with default initialized parameters.
Figure 2: The average of #Iteration and cost $C_{(x)}(\bm{\theta}^{*}_{(x)})$. Each subplot represents one qubit configuration. The blue curves indicate the cost values obtained after optimization, while the red curves show the corresponding average of #Iteration consumed to reach those cost values. Solid lines denote results with warm-starting, whereas dashed lines represent the baseline without warm-starting.
@@ IV-C Organizing targets as a multi-level tree
${C}^{(1)},{C}^{(2)},\dots,{C}^{(D)}.$ (13)
Each $\mathcal{C}^{(d)}=\{x^{(d)},C^{(d-1)}_{1},C^{(d-1)}_{2},\dots,C^{(d-1)}_{|d|}\}$ include a centroid $x^{(d)}$ and the set of tree at lower level d1d-1. Conversely, trees at level $d$ are formed by merging clusters from level d1d-1. The lower levels capture finer-grained clusters and higher levels capture more abstract groupings. This structure allows analysis at multiple scales, facilitating both local and global interpretations of the data.
Figure 1 (b) presents the 2-level tree system where level 0 has two sub-trees. The knowledge can be transferred from the centroid to sub-trees, for example: knowledge from the $\bm{\theta}^{*}_{d=0}$ is transferred to $\bm{\theta}_{d=1}$, knowledge from the $\bm{\theta}^{*}_{d=1}$ is transferred to $\bm{\theta}_{d=2}$, and so on. The optimization process on tree can be imagined as Figure 1 (c), instead of starting from $\bm{\theta}^{(0)}$, $\bm{\theta}^{*}_{d=1}$ and $\bm{\theta}^{*}_{d=2}$ can be obtained from $\bm{\theta}^{*}_{d=0}$ and $\bm{\theta}^{*}_{d=1}$, respectively. For example, we could take a hundred iterations for achieving $\bm{\theta}^{*}_{d=1}$ but only a few iterations if we start from the nearest neighbor $\bm{\theta}^{*}_{d=0}$ instead of $\bm{\theta}^{(0)}$.
@@ IV-D Deep learning method
$U(\bm{\theta}_{k})=\bigotimes_{j=0}\left(U_{\{j\}}\left(\bm{\theta}_{k\{j\}}\right)\right).$ (14)
$\bm{\theta}_{k\{j\}}$ is the parameter $k$ at layer $j$; we can share mid-layer parameter for $\bm{\theta}_{k+1}$. Meanwhile, meta‑learning methods enable fast adaptation: a meta‑learner network is trained across a distribution of $k^{\prime\prime}\ll K$ targets, producing initial PQC parameters that can be fine‑tuned in just a few gradient steps for each new target. Studies suggest that such meta‑strategies significantly improve robustness under noise [30, 31]. Together, multi-task learning and meta‑learning comprise powerful tools for scalable, sample‑efficient optimization in quantum systems deploying multi‑target optimization.
@@ V Experiments
$U(\bm{\theta})=ZXZ\prod_{i=0}^{n-1}CR_{Y}(\theta)_{i,\ (i+1)\%n}.$ (15)
The number of parameters and circuit depth of the above $n$-qubit ansatz are $3nL$ and $6L$, respectively. The problem $P$ is implemented by Python 3.10 and Pennylane 0.41.0, and the scalability is investigated by increasing $n$ and $L$ ($L=n$). The maximum #Iteration and $\tau$ are fixed at 200 and $10^{-3}$ for all cases, respectively. Higher number of qubits, harder to optimize; in other words, $P$ will require more iterations to achieve convergence. The dataset is generated quite simply; we randomly choose $k$ $n$-qubit quantum states, then sequentially optimizing each of them with $P(\bm{\theta}^{(0)},x)$ as defined in Algorithm 1. The final cost values $\mathcal{C}$ conducted by our proposed methods and baseline method are compared in the following section.
@@ V-A Effectiveness of transfer-based strategies
TABLE I: Comparison of warm-starting method and parameter estimator versus the no acceleration technique.
@@
$n$ Cost (Proposed) Cost Avg.#Iter (Proposed) Avg.#Iter
[…]
@@
$n$ Cost (Proposed) Cost Avg.#Iter (Proposed) Avg.#Iter
[…]
Table I summarizes all comparison results between transfer-based strategies and non-transfer-based strategies approaches, for two distinct methods: the warm-starting and the parameter estimator method. For each setting, we report the average of minimum cost values and the average number of iterations (Avg.#Iter) required to reach convergence. The warm-starting method is benchmarked as follows: given sets $A$ and $B$ where $|A|\gg|B|$, for each $y\in B$, we optimize $y$ with and without using knowledge from $A$. On the other hand, for the parameter estimator, we start with $B$ and an empty set $A$; for each optimized target, we move it from $B$ to $A$ until $B$ becomes an empty set. The first target will be optimized without the parameter estimator, and the latter will be transferred from $5$ nearest targets.
@@
Figure 2 shows the detailed result from the warm-starting method from 2-qubit to 5-qubit. Across all $K$, our method consistently achieves lower or comparable cost values using fewer iterations. In the 2-qubit case, the cost with warm-starting converges rapidly to a stable lower value, while the required #Iterations is significantly reduced compared to the baseline. This benefit is particularly pronounced for smaller systems, where the prior knowledge transferred from related tasks yields substantial optimization efficiency. In the 3-qubit and 4-qubit cases, the #Iteration for the warm-starting is again consistently lower, while both approaches maintain similar cost levels. For 5-qubit, as system complexity increases, the benefits of warm-starting become clearer, where both cost and #Iteration are better. However, even in these cases, warm-starting demonstrates a trend toward requiring fewer iterations to reach comparable or slightly better performance, indicating its scalability. In higher-dimensional systems (6 to 8 qubits, Figure 3, right panel), the advantage of warm-starting becomes more pronounced. The dashed lines lie consistently below the solid lines, indicating lower final cost values even in more complex quantum systems. Furthermore, the cost values obtained via transfer are relatively stable as $K$ increases, in contrast to the baseline, which exhibits higher variability. This highlights the robustness of warm-starting in high-dimensional optimization landscapes.